Shift Summary for Day shift, 06/29/99

Sydney A. Taegar (taegar@physics.arizona.edu)
Tue, 29 Jun 1999 14:58:37 -0700 (MST)

Shift Summary for Day shift, 06/29/99

On Shift: Jianbo Wang, Sydney Taegar

Hours of beam delivered at all intensities: 8.5 hours
Hours of beam delivered at nominal intensity: 8.5 hours
Hours of E832 data (special runs not included): 7.8 hours
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:

Making access to swap ET controller
Loading new tapes

========================================================================
Runs ended this shift:
Run 13757 - 4.7M events - BEAM_832 @ 8E12
Run 13758 - 0.4M events - BEAM_832 @ 8E12 (swapped ET Top East controller
during this run)
Run 13759 - 1.8M events - BEAM_832 @ 8E12 (new CSQ6 constants
and PMT gain drift ref
after this run)
Run 13760 - in progress - BEAM_832 @ 6.5E12

Tapes brought to FCC this shift:
QKK 058-060
QKC 058-060
QKS 058-060
QKB 020

========================================================================
Accesses during shift:
ST, JBW make access to swap ET controller

========================================================================
Overview of shift:

E832 data taking at 8E12 for first 3.5 hours, at 6.5E12 for remainder
of shift.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13757 7:00 BEAM_832 run at 8E12 in progress.

85 7:20 ET inefficient stripe. Reload ET.

7:29 We are sometimes dead at the beginning
of spill due to one L3 process on ktev1
which takes twice as much time per event
as other processes. Stop run 13757.

13758 7:35 Start BEAM_832 run 13758 with L3 process
brute force init in order to restart L3
processes and fix above problem.

3 ET inefficient stripe. Reload ET.

4 ET inefficient stripe. Reload ET.

5 ET inefficient stripe. Reload ET.

8 ET inefficient stripe. Reload ET.

7:57 Jianbo and Sydney go on access to replace
ET Top East controller.

8:20 Out of access. Stop run 13758.

13759 8:29 Start BEAM_832 run 13579.
DC online hists don't come up. Do fix_disp.

9:28 Mount QKN005 -> ktev2_tps3d2
QKNL05 -> ktev4_tps3d2 for onsplit.

57 9:44 KTEVANA errors (unknown crate in stream 5).
KFREND.
Jason Hamm notes that this is due
to crate 53 and would like to replace the DYC
during the next signicant downtime.

10:09 Tapes fill up, run 13759 ends.
We ask MCR for 6.5E12 for the next run.
EB installs new CSQ6 constants and updates
reference for PMT gain drift.

13760 10:25 Start BEAM_832 run 13760 at 6.5E12 writing
to new tapes QK*061-063, QKB021.

10:45 Vaulted QK*058-060, QKB020

12:00 Did gas checklist.

111 12:55 KTEVANA errors (unknown crate in stream 5).
KFREND.

143 13:38 Cr53 DYC busy. KFREND.

167 14:11 KTEVANA errors (missing FFFF word in stream 5).

15:30 Hand off to evening shift.

========================================================================
Follow up Notes on Previous Shift

ET inefficient stripes were due to corrupted L1 signal on ET
Top East controller. Swapping in a new controller has fixed
this problem.

Hogan noted that drift chamber current spikes before beginning
of spill are due to chamber ramp-up current, not current draw
from the wires. He has updated the chamber current monitor
scope to make the distinction clear to shiftees.

========================================================================
Items to Follow up on Next Shift

CsI cosmic counters U1-4 and D11-16 have no hits in the monitor plot.
Rick K suspects this may be due to incorrect pedestals. Will investigate
on next down time.

Keep running BEAM_832 at 6.5E12 until 18:00 or technician shows up
to replace tape drive tps2d2 on ktev1. At that point start
a new run at 5E12.